# A Gaussian Process View on Observation Noise and Initialization in Wide Neural Networks

### v2

@@ G.1 Geometric Decay of the regularized gradient and closeness of parameters to their initial value
$\forall\lVert x\rVert_{2}\leq 1:\lVert J(x,\theta_{t})-J(x,\theta_{0})\rVert_{2}\leq\frac{1}{\sqrt{n}}K^{\prime}C,$ (114)
$\lVert J(\theta_{t})-J(\theta_{0})\rVert_{2}\leq\frac{1}{\sqrt{n}}KC.$ (115)
@@
$2\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t-1})^{\top}\left(\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t})-\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t-1})\right)$ $=-2\int_{0}^{\eta_{0}}\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t-1})\left(\nabla_{\theta}^{2}\mathcal{L}^{\beta}\left(\theta_{t-1}-u\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t-1})\right)\right)\cdot\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t-1})du$ (124)
@@
$\leq\frac{1}{\sqrt{n}}KR_{0}+\lambda_{\max}\left(J(\theta)^{\top}J(\theta)\right)+\beta$ (127)
$\leq\frac{1}{\sqrt{n}}KR_{0}+2\lambda_{\max}(\mathbf{\Theta})+\beta$ (128)
@@ G.2 Closeness to the linearized network along the regularized gradient descent
$\sup_{t\geq 0}\lVert\theta_{t}^{\operatorname*{lin}}-\theta_{t}\rVert_{2}\leq C_{1}\frac{1}{\sqrt{n}},\quad\forall\lVert x\|_{2}\leq 1:\sup_{t\geq 0}\lVert f^{\operatorname*{lin}}(x,\theta^{\operatorname*{lin}}_{t})-f(x,\theta_{t})\rVert_{2}\leq C_{2}\frac{1}{\sqrt{n}}.$ (137)
@@
$\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta_{t})\rVert_{2}\leq\eta_{0}\frac{1}{\sqrt{n}}K^{\prime}CKR_{0}\sum_{u=0}^{t-1}\left(1-\eta_{0}c_{\beta}\right)^{u}.$ (139)
@@
$\lVert f(x,\theta_{t+1})-f^{\operatorname*{lin}}(x,\theta_{t+1})\rVert_{2}=\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta_{t})\rVert_{2}+\lVert f(x,\theta_{t+1})-f(x,\theta_{t})-\left(f^{\operatorname*{lin}}(x,\theta_{t+1})-f^{\operatorname*{lin}}(x,\theta_{t})\right)\rVert_{2}.$ (140)
@@
$\leq$ $\eta_{0}\frac{1}{\sqrt{n}}K^{\prime}CKR_{0}\left(1-\eta_{0}c_{\beta}\right)^{t}.$ (144)
@@
$\lVert f(x,\theta_{t+1})-f^{\operatorname*{lin}}(x,\theta_{t+1})\rVert_{2}\leq\eta_{0}\frac{1}{\sqrt{n}}K^{\prime}CKR_{0}\left(\sum_{u=0}^{t-1}\left(1-\eta_{0}c_{\beta}\right)^{u}+\left(1-\eta_{0}c_{\beta}\right)^{t}\right).$ (145)
@@
$\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta_{t})\rVert_{2}$ $\leq\eta_{0}\frac{1}{\sqrt{n}}K^{\prime}CKR_{0}\sum_{u=0}^{t-1}\left(1-\eta_{0}c_{\beta}\right)^{u}$ (146)
$<\eta_{0}\frac{1}{\sqrt{n}}K^{\prime}CKR_{0}\frac{1}{\eta_{0}c_{\beta}}$ (147)
$=\frac{1}{\sqrt{n}}K^{\prime}C^{2}.$ (148)
@@ Appendix I The Output of the Linearized Network is Gaussian over Random Initializations
$f(\mathbf{x}^{\prime},\theta_{0})$ and $f(\mathbf{x},\theta_{0})$ jointly converge to a Gaussian with mean zero and covariance matrix given through the NNGP-kernel $\mathbf{K}$. From this, it directly follows that $f(\mathbf{x^{\prime}},\theta_{\infty})$ converges to a Gaussian with the given mean and covariance matrices. [^14] The covariance matrix of $X+AY$, where $X$ and $Y$ are jointly Gaussian, is given by $\Sigma_{X}+A\Sigma_{Y}A^{\top}+A\Sigma_{X,Y}+\Sigma_{Y,X}A^{\top}$.

### v3

@@ G.1 Geometric Decay of the regularized gradient and closeness of parameters to their initial value
$\forall\lVert x\rVert_{2}\leq 1:\lVert J(x,\theta_{t})-J(x,\theta_{0})\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}C,$ (114)
$\lVert J(\theta_{t})-J(\theta_{0})\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}KC.$ (115)
@@
$2\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t-1})^{\top}\left(\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t})-\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t-1})\right)$ (124)
$=$ $-2\int_{0}^{\eta_{0}}\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t-1})\left(\nabla_{\theta}^{2}\mathcal{L}^{\beta}\left(\theta_{t-1}-u\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t-1})\right)\right)\cdot\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t-1})du$ (125)
@@
$\leq\frac{(\log n)^{c}}{\sqrt{n}}KR_{0}+\lambda_{\max}\left(J(\theta)^{\top}J(\theta)\right)+\beta$ (128)
$\leq\frac{(\log n)^{c}}{\sqrt{n}}KR_{0}+2\lambda_{\max}(\mathbf{\Theta})+\beta$ (129)
@@ G.2 Closeness to the linearized network along the regularized gradient descent
$\sup_{t\geq 0}\lVert\theta_{t}^{\operatorname*{lin}}-\theta_{t}\rVert_{2}\leq C_{1}\frac{(\log n)^{c}}{\sqrt{n}},\quad\forall\lVert x\|_{2}\leq 1:\sup_{t\geq 0}\lVert f^{\operatorname*{lin}}(x,\theta^{\operatorname*{lin}}_{t})-f(x,\theta_{t})\rVert_{2}\leq C_{2}\frac{(\log n)^{c}}{\sqrt{n}}.$ (138)
@@
$\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta_{t})\rVert_{2}\leq\eta_{0}\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}CKR_{0}\sum_{u=0}^{t-1}\left(1-\eta_{0}c_{\beta}\right)^{u}.$ (140)
@@
$\lVert f(x,\theta_{t+1})-f^{\operatorname*{lin}}(x,\theta_{t+1})\rVert_{2}$ (141)
$\leq$ $\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta_{t})\rVert_{2}+\lVert f(x,\theta_{t+1})-f(x,\theta_{t})-\left(f^{\operatorname*{lin}}(x,\theta_{t+1})-f^{\operatorname*{lin}}(x,\theta_{t})\right)\rVert_{2}.$ (142)
@@
$\leq$ $\eta_{0}\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}CKR_{0}\left(1-\eta_{0}c_{\beta}\right)^{t}.$ (146)
@@
$\lVert f(x,\theta_{t+1})-f^{\operatorname*{lin}}(x,\theta_{t+1})\rVert_{2}\leq\eta_{0}\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}CKR_{0}\left(\sum_{u=0}^{t-1}\left(1-\eta_{0}c_{\beta}\right)^{u}+\left(1-\eta_{0}c_{\beta}\right)^{t}\right).$ (147)
@@
$\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta_{t})\rVert_{2}$ $\leq\eta_{0}\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}CKR_{0}\sum_{u=0}^{t-1}\left(1-\eta_{0}c_{\beta}\right)^{u}$ (148)
$<\eta_{0}\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}CKR_{0}\frac{1}{\eta_{0}c_{\beta}}$ (149)
$=\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}C^{2}.$ (150)
@@ Appendix I The Output of the Linearized Network is Gaussian over Random Initializations
$f(\mathbf{x}^{\prime},\theta_{0})$ and $f(\mathbf{x},\theta_{0})$ jointly converge to a Gaussian with mean zero and covariance matrix given through the NNGP-kernel $\mathbf{K}$. From this, it directly follows that $f(\mathbf{x^{\prime}},\theta_{\infty})$ converges to a Gaussian with the given mean and covariance matrices [^14] The covariance matrix of $X+AY$, where $X$ and $Y$ are jointly Gaussian, is given by $\Sigma_{X}+A\Sigma_{Y}A^{\top}+A\Sigma_{X,Y}+\Sigma_{Y,X}A^{\top}$.. ∎